A single lane street 12ft wide goes through a semicircular tunnel with radius 10ft. How high is the tunnel at the edge of each lane?


1
Expert's answer
2020-11-16T19:20:06-0500

Let "r=10" ft, "l=12" ft.

"h=\sqrt{r^2-(\frac l2)^2}=\sqrt{10^2-6^2}=8" ft.